# Automorphism groups of affine varieties and their Lie algebras

### v1

@@ 5. Algebraic subgroups and algebraic elements
It is easily seen that $g$ is algebraic if and only if it defines a locally finite endomorphism of $\mathcal{O}(X)$. In particular, any algebraic element $g$ has a unique additive Jordan decomposition $g=g_{\mathrm{s}}+g_{\mathrm{n}}$.
@@
Let $g,h\in\operatorname{Aut}(X)$ be two algebraic elements. Then either $g,h$ are contained in an algebraic subgroup of $\operatorname{Aut}(X)$, or the group generated by $g$ and $h$ contains non-algebraic elements.
Example 5.5 (see [KrZa22, Theorem 6.4.3]).
Let $F=\langle g,h\rangle\subseteq\operatorname{Aut}({\mathbb{A}}^{2})$ be the subgroup generated by two unipotent algebraic elements
@@
Let $\operatorname{Aut}(X)$ be infinite dimensional (resp., has a nontrivial connected component). Then $\operatorname{Aut}(X)$ contains a subgroup isomorphic to either $\mathbb{G}_{\mathrm{m}}$ or $\mathbb{G}_{\mathrm{a}}$, or in other words, $\operatorname{Aut}(X)$ contains a non-torsion algebraic element.
Due to Theorem 8.4 below, Conjecture 5.6 holds for affine surfaces. It holds also provided $\operatorname{Aut}(X)$ contains a nontrivial connected set of commuting elements. Indeed, we have the following facts.
@@
Question 5.9.
Assume that the group $\operatorname{Aut}(X)$ has no nonunit algebraic element. Is it then discrete?
Notice that by Theorem 5.7 under the latter assumption every abelian subgroup of $\operatorname{Aut}(X)$ is discrete and has no torsion.
@@ 6. Nilpotency and solvability
Theorem 6.1 ([KrZa22, Theorem B]).
Let $G\subseteq\operatorname{Aut}(X)$ be an algebraically generated subgroup. If $G$ is solvable, then the following hold.
@@
Theorem 6.2 ([KrZa22, Theorem C]).
@@
where $a_{i}\in{\mathbb{K}}^{*}$ and $f_{i}\in{\mathbb{K}}[x_{1},\ldots,x_{i-1}]$. The unipotent radical $\mathcal{J}_{n}=R_{\mathrm{u}}({\rm Jonq}_{n})$ is a closed nested unipotent subgroup of $\operatorname{Aut}({\mathbb{A}}^{n})$. For $n\geq 2$ its Lie algebra $\operatorname{Lie}(\mathcal{J}_{n})$ is not nilpotent. However, the Lie algebra of a nilpotent ind-group also is nilpotent, see [KrZa22, Lemma 5.1.4(3)].
@@
It is easily seen that the $d$th member $(L_{d})_{d}$ of the lower central series of $L_{d}$ does not vanish, and so $(\operatorname{Lie}(\mathcal{J}_{2}))_{d}\neq 0$ for any $d\geq 1$.
@@
In the case of a nilpotent (resp., a solvable) subgroup $G$ one can proceed by induction on the nilpotency class of $G$ (resp., on the derived length of $G$), the case of $n(G)=1$, that is, $G$ abelian, being fixed by Theorem 5.7. In the general case, the induction works provided the following conjecture is true.
@@
If $G$ is a-generated and nilpotent (resp., solvable) then any member of its lower central series (resp., derived series) is as well a-generated.
If $G/z(G)$ is abelian, then Conjecture 6.6 holds if the following is true:
@@
Consider an admissible filtration $\operatorname{Aut}^{0}(X)=\lim A_{n}$ by irreducible affine subvarieties $A_{n}$. Let $G$ be a (nilpotent, resp., solvable) subgroup of $\operatorname{Aut}^{0}(X)$ such that $G=\langle G\cap A_{n}\rangle$ for some $n$. Then $z(G)=\langle z(G)\cap A_{m}\rangle$ for some $m$ (resp., $(G,G)=\langle(G,G)\cap A_{m}\rangle$ for some $m$), where $z(G)$ stands for the center of $G$.
For the following question, see [KrZa22, Sect. 5.2].
Question 6.9.
Let $G\subseteq\operatorname{Aut}(X)$ be a solvable or nilpotent connected subgroup. Does there exists a closed embedding $X\hookrightarrow{\mathbb{A}}^{n}$ such that $G$ extends to a subgroup of the de Jonquières group ${\rm Jonq}_{n}$? Is this true if $G$ is nested?
@@
Let $A\subseteq\operatorname{Aut}(X)$ be an irreducible algebraic subvariety which contains $\operatorname{id}_{X}$, and let $\langle A\rangle$ be the subgroup of $\operatorname{Aut}(X)$ generated by $A$. The notation $L_{A}\subseteq\operatorname{Vec}(X)$ have the same meaning as in subsection 2.1.
Question 6.11.
We wonder if the following equivalences hold:
@@
    $L_{A}\subseteq\operatorname{LND}(X)\Leftrightarrow\operatorname{Ad}(A)\subseteq\operatorname{Aut}_{Lie}(\operatorname{Vec}(X))$ is ULN $\Leftrightarrow\langle A\rangle$ is a unipotent algebraic group;
@@
    $A$ is metabelian $\Leftrightarrow\langle\operatorname{Ad}(A)\rangle_{Lie}$ is commutative;
@@
    $\langle A\rangle$ is solvable $\Leftrightarrow\langle\operatorname{Ad}(A)\rangle_{Lie}$ is;
@@
    $\langle A\rangle$ is nilpotent $\Leftrightarrow\langle\operatorname{Ad}(A)\rangle_{Lie}$ is.
@@ 7.1. $\operatorname{Aut}$-quasihomogeneous affine varieties
We say that a variety $X$ is $\operatorname{Aut}$-homogeneous (resp. $\operatorname{Aut}$-quasihomogeneous) if $\operatorname{Aut}(X)$ acts transitively on $X$ (resp. acts on $X$ with an open orbit).
This notion and the following characterizations of $\operatorname{Aut}$-quasihomogeneity for smooth affine surfaces are due to Gizatullin [Giz71b]. This was extended to normal affine surfaces by Dubouloz [Dub04].
@@
  • $X\cong\bar{X}\setminus D$ where $\bar{X}$ is a smooth projective surface and $D\subseteq{\rm reg}(\bar{X})$ is a reduced divisor with simple normal crossings, which consists of a chain of smooth rational curves.
Notice that many, but not all, smooth Gizatullin surfaces are $\operatorname{Aut}$-homogeneous. Among $\operatorname{Aut}$-homogeneous normal affine surfaces, homogeneous spaces of algebraic groups constitute a narrow subset; these were completely classified by Gizatullin [Giz71a] and Popov [Pop73], see also [KPZ17, Theorem 4.8].
@@ 7.2. Flexible affine varieties
Flexible affine varieties were introduced in [${}^{+}$13a], see also [KZ99] and [AKZ12]. For overviews on the subject see [Arz23], [${}^{+}$13b], and [CPPZ21, Section 7].
@@
Let $X$ be an affine variety. We say that a point $x\in{\rm reg}(X)$ is flexible if the tangent space $T_{x}X$ is spanned by the tangent vectors to the orbits $Ux$ of $\operatorname{{\mathbb{G}}_{\mathrm{a}}}$-subgroups $U$ of $\operatorname{Aut}(X)$. The variety $X$ is called flexible if every point $x\in{\rm reg}(X)$ is. Clearly, $X$ is flexible if one point $x\in{\rm reg}(X)$ is and the group $\operatorname{Aut}(X)$ acts transitively on ${\rm reg}(X)$.
We let $\operatorname{\mathfrak{U}}(X)\subseteq\operatorname{Aut}^{0}(X)$ be the subgroup generated by all $\operatorname{{\mathbb{G}}_{\mathrm{a}}}$-subgroups of $\operatorname{Aut}(X)$, in other words, by all unipotent elements of $\operatorname{Aut}(X)$  [^2] The subgroup $\operatorname{\mathfrak{U}}(X)$ is denoted by ${\rm SAut}(X)$ in [${}^{+}$13a]. We prefer the notation $\operatorname{\mathfrak{U}}(X)$ in order to keep ${\rm SAut}({\mathbb{A}}^{n})$ for the subgroup of $\operatorname{Aut}({\mathbb{A}}^{n})$ of volume-preserving automorphisms.. We also let $\operatorname{Aut}_{\rm alg}(X)$ be the subgroup generated by all algebraic elements of $\operatorname{Aut}(X)$. Notice that $\operatorname{Aut}_{\rm alg}(X)$ is not necessarily connected, and every $a\in\operatorname{Aut}_{\rm alg}(X)\setminus\operatorname{Aut}^{0}(X)$ is a torsion element.
The following theorem (conjectured in [AKZ12, Sect. 4.2]) was proven in [${}^{+}$13a] for affine varieties and extended in [APS14, Theorem 2] and [FKZ16, Theorem 2.12] to quasiaffine varieties.
@@
Let $X$ be a flexible smooth quasiaffine variety of dimension $\geq 2$ and $Y\subseteq X$ be a closed subset of codimension $\geq 2$. Then $X\setminus Y$ is flexible.
As an application of Theorem 7.4, let us mention the following two interpolation results. The next theorem follows immediately from [${}^{+}$13a, Theorem 4.14 and Remark 4.15].
@@
See also [${}^{+}$13a, Theorem 1.15] for an analog of Kleiman’s Transversality theorem for flexible varieties.
@@ 7.3. Flexibility of toric varieties
Let ${\mathbb{T}}=(\operatorname{{\mathbb{G}}_{\mathrm{m}}})^{n}$ stand for an algebraic $n$-torus. Recall that a ${\mathbb{T}}$-variety $X$ of dimension $n$ is a toric variety if ${\mathbb{T}}$ acts effectively on $X$ with an open orbit. Any toric affine variety $X$ of dimension $n\geq 2$ with no torus factor can be defined by a strictly convex polyhedral cone in the integer lattice ${\mathbb{Z}}^{n}$. For instance, the affine plane ${\mathbb{A}}^{2}$ with the standard diagonal torus action is defined by the first quadrant ({$({\mathbb{N}}\cup\{0\})^{2}$ of ${\mathbb{Z}}^{2}$.
@@
Such subgroups, or rather the associated locally nilpotent derivations of $\mathcal{O}(X)$, correspond to certain lattice points of ${\mathbb{Z}}^{n}$ called Demazure roots. In particular, if $X={\mathbb{A}}^{2}$ then the Demazure roots are the lattice points $(-1,k)$ and $(l,-1)$ with $k,l\geq 0$. They correspond to the locally nilpotent derivations $y^{k}\partial/\partial x$ and $x^{l}\partial/\partial y$, respectively. It is known that any normal toric variety of dimension $\geq 2$ with no torus factor is flexible, see [AKZ12]. The following theorem says more.
@@
Let $X$ be a toric affine variety of dimension at least 2. If $X$ has no torus factor and is smooth in codimention $2$, then one can find a finite collection of Demazure root subgroups such that the group generated by these acts infinitely transitively on the smooth locus ${\rm reg}(X)$.

### v2

@@ 5. Algebraic subgroups and algebraic elements
Clearly, $g$ is algebraic iff the closure $\overline{<g>}$ is an algebraic group. If $g$ is algebraic, then this abelian group has the form $(\operatorname{{\mathbb{G}}_{\mathrm{m}}})^{s}(\operatorname{{\mathbb{G}}_{\mathrm{a}}})^{t}F$ with $s\geq 0$, $t\in\{0,1\}$ and $F$ being a finite cyclic group. Every subgroup of this form in $\mathfrak{G}$ is equal to $\overline{<g>}$ for some $g\in\mathfrak{G}$.
It is easily seen that $g\in\operatorname{Aut}(X)$ is algebraic if and only if it defines a locally finite automorphism of $\mathcal{O}(X)$. In particular, any algebraic element $g$ has a unique multiplicative Jordan decomposition $g=g_{\mathrm{s}}g_{\mathrm{u}}$, where $g_{\mathrm{s}},g_{\mathrm{u}}\in\operatorname{Aut}(X)$ commute, $g_{\mathrm{s}}$ is semisimple and $g_{\mathrm{u}}$ is unipotent, see [FuKr18, Section 9.1].
@@
Let $g,h\in\operatorname{Aut}(X)$ be algebraic elements. Then either $g,h$ are contained in an algebraic subgroup of $\operatorname{Aut}(X)$, or the group generated by $g$ and $h$ contains non-algebraic elements.
The following corollary of [LPU23, Theorem 5.1] confirms the latter conjecture in dimension two.
Theorem 5.5.
Let $X$ be an affine surface and $G$ be a finitely generated subgroup of $\operatorname{Aut}(X)$. If $G$ consists of algebraic elements, then $G$ is contained in an algebraic subgroup of $\operatorname{Aut}(X)$.
Cf. also Theorem 8.7 and Corollary 8.8 below.
Due to the next results (see [BPZ24, Theorem 1.1 and Corollaries 1.2-1.3]), the Burnside problem has a positive solution in $\operatorname{Aut}(X)$.
Theorem 5.6.
Let $k$ be any field of zero characteristic and $X$ be an affine variety over $k$. If $G$ is a torsion subgroup of $\operatorname{Aut}(X)$, then any finitely generated subgroup of $G$ is finite. If $X={\mathbb{A}}^{n}_{k}$, then $G$ contains an abelian normal subgroup of finite index.
The proof uses the following result of Bass and Lubotzky [BL83, Corollary 1.2].
Theorem 5.7.
Let $k$ be an arbitrary field and $X$ be a scheme of finite type over $k$. If $G\subseteq\operatorname{Aut}(X)$ is a finitely generated subgroup, then $G$ is residually finite, i.e. the intersection of all finite index subgroups of $G$ is trivial. If $X$ is flat over ${\mathbb{Z}}$, then $G$ is virtually torsion free, in particular, the orders of finite subgroups of $G$ are bounded above.
Example 5.8 (see [KrZa24, Theorem 6.4.2]).
Let $F=\langle g,h\rangle\subseteq\operatorname{Aut}({\mathbb{A}}^{2})$ be the subgroup generated by algebraic elements
@@
Assume that $\operatorname{Aut}^{0}(X)\neq\{\operatorname{id}_{X}\}$. Then $\operatorname{Aut}^{0}(X)$ contains a subgroup isomorphic to either $\mathbb{G}_{\mathrm{m}}$ or $\mathbb{G}_{\mathrm{a}}$, or in other words, $\operatorname{Aut}(X)$ contains a non-torsion algebraic element.
Due to Theorem 8.4 below, Conjecture 5.9 holds for affine surfaces. It holds also provided $\operatorname{Aut}(X)$ contains a nontrivial connected set of commuting elements. Indeed, we have the following results.
@@
Question 5.12.
Assume that the group $\operatorname{Aut}(X)$ does not contain algebraic elements different from $\operatorname{id}_{X}$. Is it then discrete?
Notice that by Theorem 5.10, if $\operatorname{Aut}(X)$ has no algebraic element different from $\operatorname{id}_{X}$, then every abelian subgroup $H$ of $\operatorname{Aut}(X)$ has no torsion and the neutral component $H^{0}$ is trivial. Cf. also [Arz18a, Problem 20].
@@ 6. Nilpotency and solvability
Theorem 6.1 ([KrZa24, Theorem B]).
Let $G\subseteq\operatorname{Aut}(X)$ be an algebraically generated subgroup, see subsection 2.1. If $G$ is solvable, then the following hold.
@@
Theorem 6.2 ([KrZa24, Theorem C]).
@@
where $a_{i}\in{\mathbb{K}}^{*}$ and $f_{i}\in{\mathbb{K}}[x_{1},\ldots,x_{i-1}]$. The unipotent radical $\mathcal{J}_{n}=R_{\mathrm{u}}({\rm Jonq}_{n})$ is a closed nested unipotent subgroup of $\operatorname{Aut}({\mathbb{A}}^{n})$. For $n\geq 2$ its Lie algebra $\operatorname{Lie}(\mathcal{J}_{n})$ is not nilpotent. However, the Lie algebra of a nilpotent ind-group is nilpotent, see [KrZa24, Lemma 5.1.4(3)].
@@
It is easily seen that the $d$th member $(L_{d})_{d}$ of the lower central series of $L_{d}$ does not vanish, and so $(\operatorname{Lie}(\mathcal{J}_{2}))_{d}\neq 0$ for any $d\geq 1$, cf. [KrZa24, Remark 5.3.3].
@@
In the case of a nilpotent (resp., a solvable) subgroup $G$ one can try to proceed by induction on the nilpotency class of $G$ (resp., on the derived length of $G$), where the case $n(G)=1$ (i.e. $G$ abelian) is settled by Theorem 5.10. In the general case, the induction works provided the following conjecture is true.
@@
If $G$ is a-generated and nilpotent (resp., solvable) then any member of its lower central series (resp., derived series) is a-generated as well.
Let $z(G)$ stands for the center of $G$. If $G/z(G)$ is abelian, then Conjecture 6.6 holds if the following is true:
@@
Consider an admissible filtration $\operatorname{Aut}^{0}(X)=\bigcup A_{n}$ by irreducible affine subvarieties $A_{n}$. Let $G$ be a (nilpotent, resp., solvable) subgroup of $\operatorname{Aut}^{0}(X)$ such that $G=\langle G\cap A_{n}\rangle$ for some $n$. Then $z(G)=\langle z(G)\cap A_{m}\rangle$ for some $m$ (resp., $(G,G)=\langle(G,G)\cap A_{m}\rangle$ for some $m$).
The following question concerns a modified version of stable triangulation, see Section 9.8.
Question 6.9 ([KrZa24, Sect. 5.2, Question 6]).
Let $G\subseteq\operatorname{Aut}(X)$ be a solvable or nilpotent connected subgroup. Does there exists a closed embedding $X\hookrightarrow{\mathbb{A}}^{n}$ such that $G$ extends to a subgroup of the de Jonquières group ${\rm Jonq}_{n}$? Is this true if $G$ is algebraic or nested?
Remark 6.10.
The answer to the latter question is affirmative in a particular setup where $X={\mathbb{A}}^{k}$ and $G\subseteq\operatorname{Aut}({\mathbb{A}}^{k})$ is a unipotent algebraic group, see [KrZa24, Remark 1.4.2(3)]. For instance, the Nagata-Bass $\operatorname{{\mathbb{G}}_{\mathrm{a}}}$-subgroup of $\operatorname{Aut}({\mathbb{A}}^{3})$ can be triangulated in ${\mathbb{A}}^{4}$ in the above sense, being non-stably triangulable in the usual sense, cf. Example 9.37.
@@
Let $A\subseteq\operatorname{Aut}(X)$ be an irreducible algebraic subvariety which contains $\operatorname{id}_{X}$, and let $\langle A\rangle$ be the subgroup of $\operatorname{Aut}(X)$ generated by $A$. The notation $L_{A}\subseteq\operatorname{Vec}(X)$ has the same meaning as in subsection 2.1. Given an ind-subgroup $\mathfrak{G}$, the adjoint representation $\operatorname{Ad}(\mathfrak{G})\colon\mathfrak{G}\to\operatorname{GL}(\operatorname{Lie}\mathfrak{G})$ is the tangent representation at the fixed point $\operatorname{id}_{X}\in\mathfrak{G}$ of the action of $\mathfrak{G}$ on itself by inner automorphisms ${\rm Int}(g)\colon h\mapsto g\cdot h\cdot g^{-1}$, see [FuKr18, Section 7.5].
Question 6.12.
Do the following equivalences hold?
@@
    $L_{A}\subseteq\operatorname{LND}(X)\Leftrightarrow\operatorname{Ad}(A)\subseteq\operatorname{End}_{Lie}(\operatorname{Vec}(X))$ is ULN $\Leftrightarrow\langle A\rangle$ is a unipotent algebraic group;
@@
    $\langle A\rangle$ is metabelian $\Leftrightarrow\langle\operatorname{Ad}(A)\rangle_{Lie}$ is commutative;
@@
    $\langle A\rangle$ is solvable $\Leftrightarrow\langle\operatorname{Ad}(A)\rangle_{Lie}$ is solvable;
@@
    $\langle A\rangle$ is nilpotent $\Leftrightarrow\langle\operatorname{Ad}(A)\rangle_{Lie}$ is nilpotent.
@@ 7.1. $\operatorname{Aut}$-quasihomogeneous affine varieties
We say that a (not necessarily affine) variety $X$ is $\operatorname{Aut}$-homogeneous (resp. $\operatorname{Aut}$-quasihomogeneous) if $\operatorname{Aut}(X)$ acts transitively on $X$ (resp. acts on $X$ with an open orbit).
The following characterizations of $\operatorname{Aut}$-quasihomogeneity for smooth affine surfaces are due to Gizatullin [Giz71b]. This was extended to normal affine surfaces by Dubouloz [Dub04].
@@
  • $X\cong\bar{X}\setminus D$ where $\bar{X}$ is a normal projective surface and $D\subseteq\bar{X}$ is a reduced divisor with simple normal crossings contained in the smooth locus ${\rm reg}(\bar{X})$ of $\bar{X}$, which consists of a chain of smooth rational curves.
Notice that many, but not all, smooth Gizatullin surfaces are $\operatorname{Aut}$-homogeneous. See Gizatullin [Giz71a] and Popov [Pop73] for a list of normal affine surfaces that are (quasi)homogeneous spaces of algebraic groups; see also [KPZ17, Theorem 4.8].
In higher dimensions there is a classification of $\operatorname{Aut}$-homogeneous affine toric varieties and complete toric varieties, see [Arz18b]; cf. also the recent preprint [ArZa23].
@@ 7.2. Flexible affine varieties
Flexible affine varieties were introduced in [AFK+13a], see also [KZ99] and [AKZ12]. For overviews on the subject and examples we address [Arz23], [AFK+13b], and [CPPZ21, Section 7].
@@
Let $X$ be a quasiaffine variety. We say that a smooth point $x\in X$ is flexible if the tangent space $T_{x}X$ is spanned by the tangent vectors to the orbits $Ux$ of $\operatorname{{\mathbb{G}}_{\mathrm{a}}}$-subgroups $U$ of $\operatorname{Aut}(X)$. The variety $X$ is called flexible if every smooth point $x\in X$ is flexible. Clearly, $X$ is flexible if at least one smooth point $x\in X$ is and the group $\operatorname{Aut}(X)$ acts transitively on the smooth locus ${\rm reg}(X)$.
We let $\operatorname{\mathfrak{U}}(X)\subseteq\operatorname{Aut}^{0}(X)$ be the subgroup generated by all $\operatorname{{\mathbb{G}}_{\mathrm{a}}}$-subgroups of $\operatorname{Aut}(X)$, in other words, by all unipotent elements of $\operatorname{Aut}(X)$  [^1] The subgroup $\operatorname{\mathfrak{U}}(X)$ is denoted by ${\rm SAut}(X)$ in [AFK+13a]. We prefer the notation $\operatorname{\mathfrak{U}}(X)$ in order to keep ${\rm SAut}({\mathbb{A}}^{n})$ for the subgroup of $\operatorname{Aut}({\mathbb{A}}^{n})$ of volume-preserving automorphisms.. We also let $\operatorname{Aut}_{\rm alg}(X)$ be the subgroup generated by all algebraic elements of $\operatorname{Aut}(X)$. Notice that $\operatorname{Aut}_{\rm alg}(X)$ is not necessarily connected, and every algebraic element $a\in\operatorname{Aut}_{\rm alg}(X)\setminus\operatorname{Aut}^{0}(X)$ is a torsion element.
The following theorem (conjectured in [AKZ12, Sect. 4.2]) was proven in [AFK+13a] for affine varieties and extended in [APS14, Theorem 2] and [FKZ16, Theorem 2.12] to quasiaffine varieties.
@@
Let $X$ be a flexible smooth quasiaffine variety of dimension $\geq 2$, and let $Y\subseteq X$ be a closed subset of codimension $\geq 2$. Then $X\setminus Y$ is flexible.
As an application of Theorem 7.4, let us mention the following two interpolation results. The first follows immediately from [AFK+13a, Theorem 4.14 and Remark 4.15].
@@
See also [AFK+13a, Theorem 1.15] for an analogue of Kleiman’s Transversality theorem for flexible varieties.
The first nontrivial examples of flexible varieties appeared in [Giz71a] and [Giz71b]; these are $\operatorname{Aut}$-homogeneous Gizatullin affine surfaces, see subsection 7.1. See [GD75], [GD77], [KPZ17] and references therein for systematic studies on the automorphism groups of Gizatullin surfaces.
Let us mention two interesting examples of flexible affine varieties of the form $X_{n,d}={\mathbb{P}}^{n}\setminus H_{d}$, where $H_{d}$ is a hypersurface of degree $d$ in ${\mathbb{P}}^{n}$ with $(n,d)\in\{(26,3),\,(55,4)\}$, see [Giz18, Proposition 3.7 and Theorem 4.3]. These are homogeneous spaces of simple linear groups of type $E_{6}$ and $E_{7}$, respectively. The cubic form that defines $H_{3}\subseteq{\mathbb{P}}^{26}$ was first written by Camille Jordan in 1870. It is related to the configuration of $27$ lines on a smooth cubic surface in ${\mathbb{P}}^{3}$. The quartic form that defines $H_{4}\subseteq{\mathbb{P}}^{55}$ is related to the $56$ exceptional curves on a Del Pezzo surface of degree 2, and to the 28 bitangent lines of a plane quartic curve.
@@ 7.3. Flexibility of toric varieties
Let ${\mathbb{T}}=(\operatorname{{\mathbb{G}}_{\mathrm{m}}})^{n}$ stand for an algebraic $n$-torus. Recall that a ${\mathbb{T}}$-variety $X$ of dimension $n$ is a toric variety if ${\mathbb{T}}$ acts effectively on $X$ with an open orbit. One says that a toric variety $X$ has a torus factor if $X\cong Y\times(\operatorname{{\mathbb{G}}_{\mathrm{m}}})^{k}$ for some toric variety $Y$ and some $k>0$, or equivalently, if there is a nonconstant invertible regular function on $X$.
Any toric affine variety $X$ of dimension $n\geq 2$ with no torus factor can be defined by a strictly convex polyhedral cone in the integer lattice ${\mathbb{Z}}^{n}$. For instance, the affine plane ${\mathbb{A}}^{2}$ with the standard diagonal torus action is defined by the first quadrant ({$({\mathbb{N}}\cup\{0\})^{2}$ of ${\mathbb{Z}}^{2}$.
@@
Such subgroups, or rather the associated locally nilpotent derivations of $\mathcal{O}(X)$, correspond to certain lattice points of ${\mathbb{Z}}^{n}$ called Demazure roots. In particular, if $X={\mathbb{A}}^{2}$ then the Demazure roots are the lattice points $(-1,k)$ and $(l,-1)$ with $k,l\geq 0$. They correspond to the locally nilpotent derivations $y^{k}\partial/\partial x$ and $x^{l}\partial/\partial y$, respectively. It is known that any normal toric affine variety of dimension $\geq 2$ with no torus factor is flexible, see [AKZ12]. The following theorem says more.
@@
Let $X$ be a toric affine variety of dimension at least 2. If $X$ has no torus factor and is smooth in codimension $2$, then one can find a finite collection of Demazure root subgroups such that the group generated by these acts infinitely transitively on the smooth locus ${\rm reg}(X)$.